OpenStudy (anonymous):

What percent of 4/5 is 1/5?

OpenStudy (anonymous):

not a problem! this is what i got... n/100 x 4/5 = 1/5 = 1/100 x 4/5 = 4/500 = 1/125 =1/125 x n = 1/5 = 1/5 / 1/125 = 1/5 x 125/1 =125/5 =25/1 =25%

jimthompson5910 (jim_thompson5910):

Another way to look at it What percent of 4 is 1? Well the answer is 25% since 25% of 4 is 1 (ie a quarter of 4 is 1)

OpenStudy (anonymous):

Yes.